Assume that X is a Poisson random variable with μ = 22. Calculate the following probabilities. (Do not round intermediate calculations. Round your final answers to 4 decimal places.) a. P(X s 12) b. P(X = 14) c. P(X> 17) d. P(17 s X s 24)
Answer #1

Poisson Formula. Suppose we conduct a Poisson experiment, in which the average number of successes within a given region is μ. Then, the Poisson probability is:

P(x; μ) = (e) (μx) / x!

where x is the actual number of successes that result from the experiment, and e is approximately equal to 2.71828.

Given mean (\mu) = 22

a) P(X < = 12) = 0.01512

b) P(X = 14) = 0.01991

c) P(X > 17) = 0.83100

d) P(17 < = X < = 24) = P(X < = 24) - P(X < 17) = 0.71172 - 0.11704 = 0.59468